Suppose a firm producing coffee provides you with the following information: Price = R100 Quantity = 150 units of coffee Cost of production = R9500 Calculate the total revenue of the firm. (1) Is the firm currently making a profit or loss? Explain.


1
Expert's answer
2022-04-24T17:25:10-0400

The total revenue of the firm is:

"TR = 100\u00d7150 = R15,000."

Total profit is:

"TP = TR - TC = 15,000 - 9,500 = R5,500."